# the Philippine Eagle, and cosmological parameter estimation

### v2

@@ 1 Introduction
Methods for accurate and robust cosmological parameter estimation are essential for interpreting increasingly precise observational data [1, 2, 3, 4, 5]. The standard $\Lambda$CDM model, while successful in describing the large-scale dynamics of the Universe [6, 7, 8, 9, 10, 11, 12, 13], relies on statistical inference techniques to extract meaningful constraints from the data. As measurements continue to improve, so too must the methods used to analyze them.
@@
In this work, we focus on an alternative class of numerical techniques—nature-inspired, metaheuristic or evolutionary algorithms (EAs)—commonly used in optimization problems. While Markov chain Monte Carlo (MCMC) methods remain the standard in cosmology and astrophysics [40, 41], EAs offer a complementary approach, capable of efficiently exploring high-dimensional, complex, and nonlinear parameter spaces. One such EA is genetic algorithm (GA), endowed with natural evolution operations, selection, crossover, and mutation [5]. GA is also widely used in cosmology in the context of model-independent cosmological reconstruction through grammatical evolution [42, 43, 44, 45]. We also introduce two EAs, the Improved Multi-Operator Differential Evolution (IMODE) and the Philippine Eagle Optimization Algorithm (PEOA), and test their ability to recover cosmological parameters, with confidence regions provided by bootstrapping on top of optimization. It is worth emphasizing that both methods have been well established in mathematical optimization and have been used in a variety of applications, in some, outperforming GA [46, 47]  [^3] We refer the interested readers to [47] for a rigorous examination of GA, IMODE and PEOA, and other methods, to a variety of problems. We also recommend [20] for a short review of the applications of EAs in cosmology and their potential role in the future of the field.. Through this work, we introduce them to cosmology.
@@
Adding a bootstrapping mechanism to estimate confidence intervals, we investigate the performance of IMODE and PEOA in cosmological parameter estimation with mock data, together with GA and MCMC as references.
@@
The remainder of this work is organized as follows. In Section 2, we provide a brief review of the standard $\Lambda$CDM model and the key observables relevant to our analysis. Section 3 discusses the generation of our mock data sets. In Section 4, we give an overview of MCMC before detailing the EAs central to this study: GA, IMODE, and PEOA. Our results are presented in Section 5, followed by a discussion in Section 6 that addresses the strengths and limitations of each method in the context of cosmological parameter estimation. Appendix A gives pseudo-algorithms for GA, IMODE, and PEOA. Our codes implementing IMODE and PEOA can be found in GitHub.
2 The $\Lambda$CDM model
The standard cosmological model, often referred to as $\Lambda$CDM, provides a physical framework for understanding the large scale structure and dynamics of the Universe. It stitches together a cosmological constant, denoted by $\Lambda$, a Cold Dark Matter (CDM) component, and the cosmological principle to explain the observed cosmic expansion and the formation of structures such as galaxies and clusters. Additionally, it integrates with general relativity to describe the behavior of the Universe at smaller scales where nonlinearities become relevant due to gravitational collapse.
Through $\Lambda$CDM, we understand that the cosmic expansion and acceleration at sufficiently large scales are governed approximately by a simple set of equations; that admit the following analytical expression for the Hubble function,
$H\left(z\right)=H_{0}\sqrt{\Omega_{m0}\left(1+z\right)^{3}+\left(1-\Omega_{m0}\right)}\,.$ (2.1)
@@ 2 The $\Lambda$CDM model
$d_{L}(z)=\dfrac{c}{H_{0}}\left(1+z\right)\int_{0}^{z}\dfrac{dz^{\prime}}{E(z^{\prime})}\,,$ (2.2)
@@
$\delta(a)=a\,_{2}F_{1}\left(\dfrac{1}{3},1,\dfrac{11}{6};a^{3}\left(1-\Omega_{m0}^{-1}\right)\right)\,,$ (2.3)
where $\,{}_{2}F_{1}\left(a,b,c;x\right)$ is the hypergeometric function, and $a$ is the scale factor, related to the redshift via $a=1/(1+z)$. A general analytical expression exists for an arbitrary constant dark energy equation of state [48]. These are the linearized solutions to the Einstein equation in a flat Friedmann-Robertson-Lemaitre-Walker background plus ad hoc elements of a constant dark energy and CDM. Remarkably, the Universe appears to be fairly consistent with the standard cosmological model.
@@ 3 Mock cosmological data
In this section, we consider a $\Lambda$CDM model with $H_{0}=70$ km s-1 Mpc-1, $\Omega_{m0}=0.3$, and $\sigma_{8}=0.8$ (amplitude of matter power spectrum, to be introduced shortly) to generate mock data sets for testing IMODE and PEOA as well as MCMC and GA.
@@
For the ‘true’ cosmology, we consider a $\Lambda$CDM model with the parameters $(h,\Omega_{m0},\sigma_{8})=(0.7,0.3,0.8)$ where $h=H_{0}/100\ {\rm km\ s}^{-1}{\rm Mpc}^{-1}$ and $\sigma_{8}$ is the average matter power spectrum on scales of $8h^{-1}{\rm Mpc}$. The statistical methods are thus tasked to infer the true cosmology with simulated data.
@@
Figure 1: Residuals, $\Delta y=y_{\rm mock}-y_{\rm true}$, in one (colored) and a hundred (gray) realisations of mock data of expansion rate, $H(Z)$, growth rate, $f\sigma_{8}(Z)$, and supernovae distance-modulus, $\mu(Z)$; generation process described in Section 3; horizontal line at $y=0$ is obtained from fiducial $\Lambda$CDM model ($\Omega_{m0}=0.3$, $H_{0}=70$ km s-1Mpc-1, $\sigma_{8}=0.8$).
The CC are standardized clocks, providing an arguably cosmological model-independent probe to the cosmic expansion. This comes from observations of age and metalicity of temporally adjacent, passively evolving galaxies, such that the expansion rate at a redshift $z_{\rm CC}$ can be approximated by $H_{\rm CC}(z)\approx-\left(\Delta z_{\rm CC}/\Delta t\right)/\left(1+z_{\rm CC}\right)$. Thus, CC at different redshifts can be regarded as a direct accounting of late-time expansion history. Our studied CC consists of redshifts $0.07\lesssim z\lesssim 1.97$ [49, 50, 51, 52, 53, 54].
@@
$\mu(z)=5\log_{10}\left(\dfrac{d_{L}\left(z\right)}{1\ {\rm Mpc}}\right)+25\,,$ (3.1)
@@
RSD measurements utilize anisotropies in a galaxy distribution due to the differences between galaxy peculiar velocities and the Hubble flow. This keys in crucial cosmological information on the growth of structure, understood through the perturbation, $\delta(z)$, and its subsequent evolution. RSD growth data is given as $f\sigma_{8}(z)$, a product of the growth rate $f(z)=d\ln\delta(z)/d\ln a(z)=-(1+z)\delta^{\prime}(z)/\delta(z)$ and the root-mean-square matter density fluctuation in a sphere of $\sim$8 Mpc, $\sigma_{8}(z)=\sigma_{8}\delta(z)/\delta(z=0)$. In the $\Lambda$CDM model, the matter density perturbation is given by Eq. (2.3); therefore, the growth rate can be expressed as [33]
$f\sigma_{8}(z)=-\sigma_{8}(1+z)\dfrac{\delta^{\prime}(z)}{\delta(z=0)}\,,$ (3.2)
where $\sigma_{8}$ is the amplitude of the matter power spectrum, setting the scale of perturbations. Note that the density fluctuation/constrast can be explicitly written in terms of the redshift, $z=a^{-1}-1$, in $\Lambda$CDM as
$\delta(z)=\dfrac{1}{1+z}\,_{2}F_{1}\left(\frac{1}{3},1,\frac{11}{6};\dfrac{\left(1-\Omega_{m0}^{-1}\right)}{(1+z)^{3}}\right)\,.$ (3.3)
This can be differentiated analytically to substitute into Eq. (3.2) to efficiently numerically evaluate $f\sigma_{8}(z)$ for comparison with simulated data.
@@ 4.1 MCMC
For our purposes, we utilize the python package emcee [61] to perform MCMC, and obtain posterior estimates of the cosmological parameters given mock data sets. The likelihoods are given as follows, based on the standard $\Lambda$CDM model (Section 2). For the expansion rate, we compare the model to the data using a likelihood ${\cal L}$ given by
$\log{\cal L}\propto-\dfrac{1}{2}\sum_{z_{\rm D}}\left(\dfrac{H_{\rm M}\left(z_{\rm D}\right)-H_{\rm D}(z_{\rm D})}{\Delta H_{\rm D}(z_{\rm D})}\right)^{2}\,,$ (4.2)
@@
$\log{\cal L}\propto-\dfrac{1}{2}\sum_{z_{\rm D}}\left(\dfrac{\mu_{\rm M}(z_{\rm D})-\mu_{\rm D}(z_{\rm D})}{\Delta\mu_{\rm D}(z_{\rm D})}\right)^{2}\,,$ (4.3)
@@
$\begin{split}\log{\cal L}\propto-\dfrac{1}{2}\sum_{z_{\rm D}}&\left(\dfrac{f\sigma_{8,{\rm M}}(z_{\rm D})-f\sigma_{8,{\rm D}}(z_{\rm D})}{\Delta f\sigma_{8,{\rm D}}(z_{\rm D})}\right)^{2}\,.\end{split}$ (4.4)
As explained previously, for the purposes of this work, our mock data sets do not take into account the full covariance of the observed data [^5] The covariance of real data can be accounted for by replacing the weights in (4.2-4.4) by the inverse of the covariance matrix..
@@ 4.2 Genetic Algorithm
The GA is part of a broader EAs, which aim to optimize the quality of a population of solutions rather than focusing solely on individual solutions. GA has found diverse applications, including in high-energy physics and gravitational wave astronomy [62, 63], and was initially introduced in cosmology as a tool to mitigate biases in model selection when analyzing dark energy [42]. Subsequent studies have demonstrated GA’s effectiveness in estimating cosmological parameter uncertainties [43, 44], selecting kernel functions in Gaussian process regression [64], optimizing neural network architectures [65], enhancing spectroscopic modeling [66, 67, 68], and aiding model selection with information criteria [69]. For a recent overview of GA in cosmology, we refer our readers to [5, 20], and for further applications, to [70, 71].
Begin Input: objective function $f$, dimension $D$, search space bounds $X_{\textrm{min}}$, $X_{\textrm{max}}$, crossover rate $cr$, archive size $m$ Set current population $k=1$, maximum function evaluations $Max\text{-}FES$, initial population size $PS$ Evaluate $f(X)$ and sort $k\leftarrow k+1$ $FES\leq Max\text{-}FES$? Choose the best $m$ individuals to form the elite pool. Apply a tournament selection with size $TC$ and fill the selection pool. Select any two individuals (parents) from the pool and generate an offspring with crossover rate $cr$ Mutate the offspring. Generate the new population by combining the elite and the pool of offsprings, maintaining the population size Sort new population Output: best solution EndYesNo
@@
We employ the built-in ga command in MATLAB, using its default settings for unconstrained problems. The population size is initially set to 50, and 5% of the population is preserved as an elite pool. For parent selection, the algorithm assigns each candidate a segment of a line proportional to its scaled fitness value and selects parents by stepping uniformly along this line. Crossover is performed by randomly combining genes from two parents based on a binary mask, with a crossover rate of 80%. Mutation is handled by perturbing each gene by adding Gaussian noise with mean zero. The standard deviation of this noise is initialized based on the population range and gradually decreases over generations.
@@ 4.3 Improved Multi-Operator Differential Evolution
Begin Input: objective function $f$, dimension $D$, search space bounds $X_{\textrm{min}}$, $X_{\textrm{max}}$ Set $G=1$, $ls=0.1$, maximum function evaluations $Max\text{-}FES$, initial population size $NP_{G}$ Evaluate $f(X)$ and sort $G\leftarrow G+1$ $FES\leq Max\text{-}FES$? Update $NP_{G}$ and assign the number of solutions $NP_{op}$ to each operator Operator 2: current-to-$\phi$best without archive Operator 1: current-to-$\phi$best with archive Operator 3: weighted-rand-to-$\phi$best Generate the new population based on the operators and update the archive Calculate $NP_{op}$ based on solutions’ quality and diversity Sort new population $FES>0.85\,Max\text{-}FES$ and $rand<ls$? Local search from the best point in current population Did the local search improve the solution? Update the best solution $ls\leftarrow 0.1$ $ls\leftarrow 0.01$ Output: best solution EndYesYesYesNoNoNo
@@
The inputs of IMODE are the objective function, dimension, and bound constraints. At the beginning of the IMODE process, an initial population from the search space is generated. The objective function using each population member is calculated and then sorted in ascending order. IMODE divides the population into several sub-populations, which are all evolved using three mutation operators. Archiving is done to preserve population diversity. The three operators are given as follows:
Operator 1: $v_{i,j}=x_{i,j}+F_{i}\times\left(x_{\phi,j}-x_{i,j}+x_{r_{1},j}-x_{r_{2},j}\right)$,
Operator $2:v_{i,j}=x_{i,j}+F_{i}\times\left(x_{\phi,j}-x_{i,j}+x_{r_{1},j}-x_{r_{3},j}\right)$,
Operator 3: $v_{i,j}=F_{i}\times\left(x_{r_{1},j}+\left(x_{\phi,j}-x_{r_{3},j}\right)\right)$,
where $r_{1},r_{2},r_{3}\neq i$ are randomly generated integers, $\vec{x}_{r_{1}},\vec{x}_{r_{3}}$ are randomly chosen from the population, $\vec{x}_{\phi}$ is selected from the top $10\%$ members of the population, and $\vec{x}_{r_{2}}$ is randomly selected from the union of the archive and the population.
[…]
@@
We take the objective function for IMODE to be the likelihoods given by Eqs. (4.2-4.4). We consider an initial population size of 54, minimum number of individuals 4, maximum number of function evaluations (Max-FES) 500, and local search probability 0.1. Max-FES fixes the minimum and maximum local search evaluations [46]. For more details, our implementation can be found in github.com/ErikaAntonette/Philippine-Eagle-Optimization-Algorithm.

### v3

# and cosmological parameter estimation

@@ 1 Introduction
[…]
@@
In this work, we primarily focus on the applicability of an alternative class of numerical techniques—nature-inspired, metaheuristic or evolutionary algorithms (EAs)—which are commonly used in optimization problems, for cosmological parameter estimation. While Markov chain Monte Carlo (MCMC) methods remain the standard in cosmology and astrophysics [40, 41], EAs offer a complementary approach, capable of efficiently exploring high-dimensional, complex, and nonlinear parameter spaces. One such EA is genetic algorithm (GA), endowed with natural evolution operations, selection, crossover, and mutation [5, 42]. GA is also widely used in cosmology in the context of model-independent cosmological reconstruction through grammatical evolution [43, 44, 45, 46]. We also introduce two EAs, the Improved Multi-Operator Differential Evolution (IMODE) and the Philippine Eagle Optimization Algorithm (PEOA), and test their ability to recover cosmological parameters, with confidence regions provided by bootstrapping on top of optimization. It is worth emphasizing that both methods have been well established in mathematical optimization and have been used in a variety of applications, in some, outperforming GA [47, 48]  [^1] We refer the interested readers to [48] for a rigorous examination of GA, IMODE and PEOA, and other methods, to a variety of problems. We also recommend [20] for a short review of the applications of EAs in cosmology and their potential role in the future of the field.. Through this work, we introduce them to cosmology.
@@
Adding a bootstrapping mechanism to estimate confidence intervals, we investigate the performance of IMODE and PEOA in cosmological parameter estimation with a oversimplified mock data (diagonal covariance), together with GA and MCMC as references. We emphasize that the uncertainties using this approach approximate frequentist intervals obtained by resampling the data, and are not Bayesian. Most importantly, the bootstrap method is a natural path to estimate errors in optimization strategies such as EAs that do not come with their own uncertainty estimates. This serves to draw a baseline for future more complicated, realistic analyses.
@@
[…]
$H\left(z\right)=H_{0}\sqrt{\Omega_{m0}\left(1+z\right)^{3}+\left(1-\Omega_{m0}\right)}\,.$ (1)
@@
$d_{L}(z)=\dfrac{c}{H_{0}}\left(1+z\right)\int_{0}^{z}\dfrac{dz^{\prime}}{E(z^{\prime})}\,$, (2)
@@
$\delta(a)=a\,_{2}F_{1}\left(\dfrac{1}{3},1,\dfrac{11}{6};a^{3}\left(1-\Omega_{m0}^{-1}\right)\right)\,$, (3)
where $_{2}F_{1}\left(a,b,c;x\right)$ is the hypergeometric function, and $a$ is the scale factor, related to the redshift via $a=1/(1+z)$. A general analytical expression exists for an arbitrary constant dark energy equation of state [49]. These are the linearized solutions to the Einstein equation in a flat Friedmann-Lemaítre-Robertson-Walker background plus ad hoc elements of a constant dark energy and CDM. Remarkably, the Universe appears to be fairly consistent with the standard cosmological model.
@@ 3 Mock cosmological data
[…]
@@
[…]
@@
[…]
The CC are standardized clocks, providing an arguably cosmological model-independent probe to the cosmic expansion. This comes from observations of age and metallicity of temporally adjacent, passively evolving galaxies, such that the expansion rate at a redshift $z_{\rm CC}$ can be approximated by $H_{\rm CC}(z)\approx-\left(\Delta z_{\rm CC}/\Delta t\right)/\left(1+z_{\rm CC}\right)$. Thus, CC at different redshifts can be regarded as a direct accounting of late-time expansion history. Our studied CC consists of redshifts $0.07\lesssim z\lesssim 1.97$ [50, 51, 52, 53, 54, 55].
@@
$\mu(z)=5\log_{10}\left(\dfrac{d_{L}\left(z\right)}{1\ {\rm Mpc}}\right)+25\,$, (4)
@@
We must emphasize that SNe data constrains only relative distances and requires an additional calibration of their absolute magnitude or an external anchor to estimate luminosity distances. On the other hand, our simplified mock data will be based on theoretically calibrated SNe since the work mainly concerns setting a baseline comparison of EAs and MCMC. However, this gives unrealistic (sub-percent) precisions for cosmological parameters, such as the Hubble constant, which we warn should not be overthought.
[…]
$f\sigma_{8}(z)=-\sigma_{8}(1+z)\dfrac{\delta^{\prime}(z)}{\delta(z=0)}\,$, (5)
[…]
$\delta(z)=\dfrac{1}{1+z}\,_{2}F_{1}\left(\frac{1}{3},1,\frac{11}{6};\dfrac{\left(1-\Omega_{m0}^{-1}\right)}{(1+z)^{3}}\right)\,.$ (6)
This can be differentiated analytically to substitute into Eq. (5) to efficiently numerically evaluate $f\sigma_{8}(z)$ for comparison with simulated data.
@@ 4.1 MCMC
[…]
$\log{\cal L}\propto-\dfrac{1}{2}\sum_{z_{\rm D}}\left(\dfrac{H_{\rm M}\left(z_{\rm D}\right)-H_{\rm D}(z_{\rm D})}{\Delta H_{\rm D}(z_{\rm D})}\right)^{2}\,$, (8)
@@
$\log{\cal L}\propto-\dfrac{1}{2}\sum_{z_{\rm D}}\left(\dfrac{\mu_{\rm M}(z_{\rm D})-\mu_{\rm D}(z_{\rm D})}{\Delta\mu_{\rm D}(z_{\rm D})}\right)^{2}\,$, (9)
@@
$\begin{split}\log{\cal L}\propto-\dfrac{1}{2}\sum_{z_{\rm D}}&\left(\dfrac{f\sigma_{8,{\rm M}}(z_{\rm D})-f\sigma_{8,{\rm D}}(z_{\rm D})}{\Delta f\sigma_{8,{\rm D}}(z_{\rm D})}\right)^{2}\,.\end{split}$ (10)
In addition to the above likelihoods (8-10), we consider the flat/uniform unit cube priors on $(h,\Omega_{m0},\sigma_{8})$, as shown in Table 4.1, to estimate the posterior using MCMC through Bayes’ theorem (7). Furthermore, we implement the sampling routine starting at a maximum of the likelihood (obtained using Nelder-Mead local search) with a total of 60,000 steps and 100 burn-in steps. Readers interested can view our implementation in GitHub.
Table 1: MCMC priors used in this work, where $h=H_{0}/\left({100\,{\rm km}\,{\rm s}^{-1}{\rm Mpc}^{-1}}\right)$ where $H_{0}$ is the Hubble constant.
Parameter Prior Values
$h$ $[0,1]$
$\Omega_{m0}$ $[0,1]$
$\sigma_{8}$ $[0,1]$
As explained previously, for the purposes of this work, our mock data sets do not take into account the full covariance of the observed data [^4] The covariance of real data can be accounted for by replacing the weights in (8-10) by the inverse of the covariance matrix..
@@ 4.2 Genetic Algorithm
Begin Input: objective function $f$, dimension $D$, search space bounds $X_{\textrm{min}}$, $X_{\textrm{max}}$, crossover rate crcr, archive size $m$ Set current population $k=1$, maximum function evaluations $Max\text{-}FES$, initial population size PSPS Evaluate $f(X)$ and sort $k\leftarrow k+1$ $FES\leq Max\text{-}FES$? Choose the best $m$ individuals to form the elite pool. Apply a tournament selection with size $TC$ and fill the selection pool. Select any two individuals (parents) from the pool and generate an offspring with crossover rate crcr Mutate the offspring. Generate the new population by combining the elite and the pool of offsprings, maintaining the population size Sort new population Output: best solution EndYesNo
@@
The GA is part of a broader EAs, which aim to optimize the quality of a population of solutions rather than focusing solely on individual solutions. GA has found diverse applications, including in high-energy physics and gravitational wave astronomy [63, 64], and was initially introduced in cosmology as a tool to mitigate biases in model selection when analyzing dark energy [43]. Subsequent studies have demonstrated GA’s effectiveness in estimating cosmological parameter uncertainties [44, 45], selecting kernel functions in Gaussian process regression [65], optimizing neural network architectures [66], enhancing spectroscopic modeling [67, 68, 69], and aiding model selection with information criteria [70]. For a recent overview of GA in cosmology, we refer our readers to [5, 20, 42], and for further applications, to [71, 72].
@@
GA, mathematically, is a global optimization method that is specifically designed to overcome shortcomings of classical search routines such as prior dependencies and tendency to become trapped in local minima [73, 74, 75, 76, 77]. Additionally, GA can be viewed as a population method; where a population $\bf p$} is acted upon by a set of probabilistic operators $\bf O$}, until a termination criterion is met, or the maximum number of generations is reached. The population is ranked according to their fitness $f$ and the random operators are selection ${\bf O}_{\rm Sel}$, crossover ${\bf O}_{\rm Cr}$, and mutation ${\bf O}_{\rm Mut}$, each of which are a function of their hyperparameters such as the mutation type and probability [78, 79]. GA mimics natural selection through the action of an operator that can be viewed macroscopically through the population or microscopically via the genetic level, e.g., in our context an individual would be a set of cosmological parameters $(H_{0},\Omega_{m0},\sigma_{8})$ and the action of an operator $(H_{0},\Omega_{m0},\sigma_{8})\rightarrow(H_{0}^{\prime},\Omega_{m0}^{\prime},\sigma_{8}^{\prime})$. Intrinsically, each operator can be thought of as applying a probability density function to an individual’s genetic make-up to keep diversity at the same time looking for possible improvements in future generations. Then, the population after the $k$th generation is [73]
${\mathbf{p}}_{k+1}(f_{k+1})={\mathbf{O}}_{\rm Mut}{\mathbf{O}}_{\rm Cr}{\mathbf{O}}_{\rm Sel}{\mathbf{p}}_{n}(f_{n})\,$, (11)
and the final evolved population can be written as a successive application of the operators:
${\mathbf{p}}_{K}(f_{K})=\left({\mathbf{O}}_{\rm Mut}{\mathbf{O}}_{\rm Cr}{\mathbf{O}}_{\rm Sel}\right)^{N}{\mathbf{p}}_{1}(f_{1})$ (12)
where ${\mathbf{p}}_{1}$ is the first (prior) population distribution. The population can be expected to cluster toward the solution in the limit when $K$ becomes increasingly large.
We employ the built-in ga command in MATLAB, using its default settings for unconstrained problems. The population size is initially set to 50, and 5% of the population is preserved as an elite pool. For parent selection, the algorithm assigns each candidate a segment of a line proportional to its scaled fitness value and selects parents by stepping uniformly along this line. Crossover is performed by randomly combining genes from two parents based on a binary mask, with a crossover rate of 80%. Mutation is handled by perturbing each gene by adding Gaussian noise with mean zero. The standard deviation of this noise is initialized based on the population range and gradually decreases over generations. Table 4.2 presents the default parameter values of the ga solver in MATLAB.
Table 2: The default values of the hyperparameters of GA that are used in the simulations.
Parameter Type/Value
dimension 3
maximum number of function evaluations 500
maximum number of generations 10
population size 50
selection stochastic uniform
crossover scattered
crossover rate 0.8
mutation Gaussian
elitism percentage 0.05
@@ 4.3 Improved Multi-Operator Differential Evolution
[…]
@@
The inputs of IMODE are the objective function, dimension, and bound constraints. At the beginning of the IMODE process, an initial population from the search space is generated. The objective function using each population member is calculated and then sorted in ascending order. IMODE divides the population into several sub-populations, which are all evolved using three mutation operators. Archiving is done to preserve population diversity. In particular, inferior solutions replaced during the evolution process are stored in the archive, which serves as an external memory. These archived individuals are later used in the mutation step to introduce new variations and prevent premature convergence. This mechanism helps the algorithm maintain exploration capability and improves its ability to escape local minima. The three operators are given as follows:
  • 1.
    Operator 1 (current-to-best with archive):
    $v_{i,j}=x_{i,j}+F_{i}\times\left(x_{\phi,j}-x_{i,j}+x_{r_{1},j}-x_{r_{2},j}\right),$
  • 2.
    Operator 2 (current-to-best without archive):
    $v_{i,j}=x_{i,j}+F_{i}\times\left(x_{\phi,j}-x_{i,j}+x_{r_{1},j}-x_{r_{3},j}\right),$
  • 3.
    Operator 3 (weighted-rand-to-best):
    $v_{i,j}=F_{i}\times\left(x_{r_{1},j}+\left(x_{\phi,j}-x_{r_{3},j}\right)\right),$
where $r_{1},r_{2},r_{3}\neq i$ are randomly generated integers, $\vec{x}_{r_{1}},\vec{x}_{r_{3}}$ are randomly chosen from the population, $\vec{x}_{\phi}$ is selected from the top $10\%$ members of the population, and $\vec{x}_{r_{2}}$ is randomly selected from the union of the archive and the population. The difference vectors $x_{r_{1},j}-x_{r_{2},j}$ and $x_{r_{1},j}-x_{r_{3},j}$ introduce stochastic perturbations that drive exploration in the population. The term $x_{r_{1},j}-x_{r_{2},j}$, where one of the indices may come from the archive, allows larger and more diverse movements that help the algorithm escape local minima and maintain population diversity. In contrast, $x_{r_{1},j}-x_{r_{3},j}$ uses only individuals from the current population, resulting in smaller, more correlated steps that promote exploitation and fine-tuning around promising regions.
The scaling factor $F_{i}$ is assigned based on success-historical memory. Operators 1 and 2 move the member in the current population to the best points with and without archiving, respectively. Operator 3 is a weighted random-member-to-best operator. The size of each sub-population $\left(NP_{op}\right$., op $\left.=1,2,3\right)$ is iteratively adjusted based on the diversity of the sub-populations and the quality of the solutions. The size of the population per generation $\left(NP_{G}\right)$ is also reduced linearly. After the mutation operators are carried out, crossover is implemented randomly to create a new set of solutions.
@@
We take the objective function for IMODE to be the likelihoods given by Eqs. (8-10). We consider an initial population size of 54, minimum number of individuals 4, maximum number of function evaluations (Max-FES) 500, and local search probability 0.1. Max-FES fixes the minimum and maximum local search evaluations [47]. For more details, our implementation can be found in github.com/ErikaAntonette/Philippine-Eagle-Optimization-Algorithm.
We use the default parameter values provided in the main reference paper of IMODE [47], which are listed in Table 4.3.
Table 3: The default values of the hyperparameters of IMODE that are used in the simulations.
Parameter Value
dimension 3
maximum number of function evaluations 500
initial population size 54
minimum population size 4
minimum number of function evaluations to begin local search 425
local search probability 0.1
archive rate 2.6
memory size 60